# Dynamical disequilibrium in dwarf galaxies: rethinking gas dynamics, rotation curves, and dark matter inference

### v1

@@ 6 Conclusions
    Disequilibrium is ubiquitous. Across our sample, both the full equilibrium model (Equation (2)) and the naïve centrifugal balance assumption (Equation (3)) frequently fail to recover the true circular velocity, with typical fractional deviations of $\geq 10$ per cent. The inner few kiloparsecs, which provide the greatest leverage for dark matter profile constraints, are also the most strongly perturbed with fractional deviations typically much greater than $10$ per cent.
@@
A second limitation is the numerical implementation of feedback. While COLIBRE incorporates a self-consistent treatment of the cold and dense phases of the ISM, and an enhanced sampling of the gravitational potential, mitigating spurious baryon–DM energy transfer, hence making it an ideal choice for analyses of gas dynamics, our results will inevitably depend on the strength and coupling of feedback, hence remain model-dependent. Our cross-checks with the hybrid AGN feedback dual versions confirm that our main results are robust to these numerical choices: despite different calibration values and injection schemes, the nature and spatial pattern of disequilibrium are qualitatively unchanged. Ultimately, verifying the generality of our findings will require similarly resolved analyses in other simulation suites. Comparisons with recent FIRE-3 work using similar diagnostics (Sands2024) further suggest that our qualitative conclusions are not simulation specific.
@@
The following software packages were used in this work: numpy (Numpy), scipy888docs.scipy.org/doc/, scikit-learn (Scikit-learn), numba (Numba), astropy (Astropy), h5py (h5py), hbt-herons (ForouharMoreno2025), soap (McGibbon2025), SwiftSimIO (swiftsimio), SWIFTGalaxy (SWIFTGalaxy).
@@ Appendix A SPH Field Sampling in a Warped Gas Disc: A Representative Example
Appendix B Classification Algorithm

### v2

@@ 6 Conclusions
    Disequilibrium is ubiquitous. Across our sample, both the full equilibrium model (Equation (2)) and the naïve centrifugal balance assumption (Equation (3)) frequently fail to recover the true circular velocity, with typical fractional deviations of $\geq 10$ per cent. The inner few kiloparsecs, which provide the greatest leverage for dark matter profile constraints, are also the most strongly perturbed with fractional deviations typically much greater than $10$ per cent. In particular, we find a population-wide tendency for the rotational velocity to systematically underestimate the true circular velocity in the inner disc, while a minority of regions exhibit the opposite bias; this bidirectional scatter can naturally result in a wide diversity of inferred inner rotation-curve shapes, including both cored and centrally concentrated profiles.
@@
A second limitation is the numerical implementation of feedback. While COLIBRE incorporates a self-consistent treatment of the cold and dense phases of the ISM, and an enhanced sampling of the gravitational potential, mitigating spurious baryon–DM energy transfer, hence making it an ideal choice for analyses of gas dynamics, our results will inevitably depend on the strength and coupling of feedback, hence remain model-dependent. Our cross-checks with the hybrid AGN feedback dual versions confirm that our main results are robust to these numerical choices: despite different calibration values and injection schemes, the nature and spatial pattern of disequilibrium are qualitatively unchanged. Ultimately, verifying the generality of our findings will require similarly resolved analyses in other simulation suites. Comparisons with recent FIRE-3 work using similar diagnostics (Sands et al., 2024) further suggest that our qualitative conclusions are not simulation specific. Related evidence for systematic departures from equilibrium and biased circular-velocity recovery has also been reported by Jahn et al. (2023) within idealised simulations using the SMUGGLE model, indicating that these effects arise across a range of numerical implementations and modelling strategies, even outside fully cosmological contexts. Within the EDGE simulation suite, Rey et al. (2024) found that inner circular velocities are recovered only in a rare subset of quiescent dwarf galaxies that temporarily host well-ordered Hi discs. In particular, their results reinforce our conclusions that rotation-curve–based mass estimates critically depend on the dynamical state of the gas – which they also found tightly linked to morphology and star formation history, and that corrections for pressure gradients do not systematically improve recovery of the underlying gravitational potential.
@@
The following software packages were used in this work: numpy (Harris et al., 2020), scipy888docs.scipy.org/doc/, scikit-learn (Pedregosa et al., 2011), numba (Lam et al., 2015), astropy (Astropy Collaboration et al., 2022), h5py (Collette et al., 2022), hbt-herons (Forouhar Moreno et al., 2025), soap (McGibbon et al., 2025), SwiftSimIO (Borrow and Borrisov, 2020), SWIFTGalaxy (Oman, 2025).
@@ Appendix A SPH Field Sampling in a Warped Gas Disc: A Representative Example
Appendix B Accounting for Disequilibrium: Sensitivity to Threshold Variations
As discussed in Sec. 4.2, the quantitative contribution of individual physical drivers to the disequilibrium budget depends to some extent on the adopted criteria used to associate gas elements as affected by stellar feedback, AGN activity, and self-gravitating clumps. While our fiducial selections are physically motivated and empirically calibrated, it is important to assess the robustness of the resulting trends to reasonable variations in these somewhat arbitrary choices.
To this end, we repeat the full analysis using two alternative sets of tagging criteria. In a permissive configuration, we loosen the spatial apertures and selection thresholds such that the total disc area associated with each driver is increased by approximately a factor of two relative to the fiducial case. Conversely, in a restrictive configuration, we tighten the criteria so that the tagged area is reduced by a comparable factor. In all cases, the same physical definitions are retained, and only the extent of the tagged regions is modified.
Table 1: Summary of the tagging criteria used to associate gas with the different disequilibrium drivers in the fiducial analysis and in the permissive and restrictive variants explored in Appendix B. The permissive and restrictive configurations are constructed to roughly double or halve, respectively, the total disc area tagged by each mechanism relative to the fiducial case, while preserving the same physical definitions.
Mechanism Parameter description Symbol [unit] Restrictive Fiducial Permissive
Stellar feedback Maximum distance from YS or from CCSNe kicked gas $R_{\rm SF}[\rm kpc]$ $\approx 0.423$ =1.000=1.000 $\approx 2.450$
Self-gravitating clumps Free-fall to orbital time ratio $t_{\rm ff}/t_{\rm orb}[\rm dimensionless]$ $\approx 0.094$ $\approx 0.111$ $\approx 0.131$
AGN feedback Maximum distance from active SMBH $R_{\rm AGN}[\rm kpc]$ $\approx 1.414$ =2.000=2.000 $\approx 2.828$
Figure 11 shows the resulting mass-weighted distributions of the disequilibrium parameter $\varepsilon$ for each driver under the three configurations, compared to the global distribution. As expected, loosening the cuts increases the overlap with the near-equilibrium regime, while tightening them preferentially selects gas at larger $|\varepsilon|$. Importantly, however, the overall shapes of the distributions are preserved to a good degree. Self-gravitating clumps preferentially populate the positive side of the distribution, while AGN feedback remains strongly associated with large negative $\varepsilon$.
Stellar feedback exhibits the largest shape variation, particularly in the transition from the fiducial to the permissive configuration: in the latter case, the distribution becomes increasingly weighted toward the near-equilibrium regime, with a reduced relative contribution at negative $\varepsilon$. This behaviour indicates that overly permissive spatial selections (e.g. circular masks of order $\approx 2.5~\rm kpc$ around young stellar particles or feedback-affected gas) begin to include substantial amounts of ambient or recovering gas. This effect naturally drives the mild differences in recovery fractions observed across configurations and directly motivates the calibration procedure adopted for the fiducial cuts.
Figure 11: Self-normalised mass-weighted distributions of the disequilibrium parameter $\varepsilon$, $\mathrm{W}_{\rm I}(\varepsilon)/\Sigma\mathrm{W}_{\rm I}$, for gas associated with each physical driver under different tagging criteria (summarised in Table 1). Each panel shows the distribution for one driver (stellar feedback, self-gravitating clumps, AGN feedback), comparing the fiducial selection (black) to permissive (orange) and restrictive (blue) variants. The grey shaded curve indicates the global gas distribution. Vertical dashed lines mark the quasi-equilibrium range, $|\varepsilon|\leq 0.2$. While the absolute normalisation and degree of overlap with the low-$|\varepsilon|$ regime vary with the tagging criteria, the qualitative association between each driver and characteristic disequilibrium regimes is robust.
The impact of these variations on the integrated disequilibrium budget is summarised in Fig. 12, which shows the fractional contribution of each driver – including overlaps – to the total gas mass tagged by the union of the three mechanisms. As anticipated, the relative ranking of the drivers is robust to changes in the tagging criteria: stellar feedback consistently dominates the budget, while AGN feedback and self-gravitating clumps contribute at a secondary but non-negligible level. The substantial overlap fractions are likewise largely preserved across configurations, reinforcing the conclusion that disequilibrium is frequently driven by multiple mechanisms acting simultaneously.
The only clear systematic difference induced by varying the thresholds is a redistribution of mass between the individual stellar feedback and AGN feedback components and their mutual overlap. This behaviour is expected, as the tagging criteria for both mechanisms are purely geometrical: increasing the characteristic radii naturally enlarges the region of intersection, reducing the fraction uniquely attributed to each driver (and vice versa for more restrictive cuts). By contrast, analogous trends are much weaker for self-gravitating clumps. This reflects the fundamentally different nature of their selection: varying the threshold in $t_{\rm ff}/t_{\rm orb}$ does not simply expand or contract contiguous regions, but can instead select spatially disconnected structures. As a result, changes in the overlap between clumps and the other drivers are less straightforward and do not scale monotonically with the strictness of the cut (in fact, they are close to independent of them). Notably, the differences discussed here remain limited to the $\approx 5$-$10$ per cent level across configurations, indicating that they do not qualitatively affect the inferred disequilibrium budget.
For completeness, and in reference to the integrated percentages discussed in the final paragraphs of Sec. 4.2, the restrictive and permissive configurations respectively account for $\approx 20$ and $50$ per cent of the gas mass in the near-equilibrium regime ($|\varepsilon|\leq 0.2$), and $\approx 70$ and $90$ per cent in the extreme disequilibrium regime ($|\varepsilon|\geq 1.0$). In the intermediate regime ($0.2<|\varepsilon|<1.0$), they encompass $\approx 35$ and $65$ per cent of the global gas mass. The systematic and differential response across $|\varepsilon|$ regimes further motivates the fiducial configuration as a balanced and optimised choice.
Taken together, these tests demonstrate that while the absolute normalisation of the tagged mass varies with the adopted thresholds, the qualitative association between physical drivers and characteristic disequilibrium regimes – and their relative partitioning – remains remarkably stable. This robustness lends confidence to the fiducial tagging strategy and supports the broader conclusion that significant departures from equilibrium in the disc arise from the combined action of stellar feedback, AGN activity, and gravitational instability, rather than from any single dominant mechanism.
Figure 12: Fractional contribution of each physical driver to the total disequilibrium budget under fiducial, permissive, and restrictive tagging criteria (see Table 1). Coloured regions indicate the contributions from stellar feedback, AGN feedback, and self-gravitating clumps, while hatched regions mark overlaps between drivers as indicated in the legend. Although the absolute fractions vary with the adopted cuts in different $\varepsilon$ regimes, the hierarchy of drivers and of overlap regions remain stable across all configurations.
Appendix C Classification Algorithm